OpenStudy (anonymous):

Find all absolute extrema of the given function on each indicated interval. f(x) = x^3 - 3x + 1 on a) [0,2] b) [-3,2]

OpenStudy (anonymous):

ok so i substitute them in?

zepdrix (zepdrix):

You want to take the derivative and find critical points. Are you able to do that portion onegirl? c:

OpenStudy (anonymous):

yes i know how to find the critical points

OpenStudy (anonymous):

or not guess this is something else

zepdrix (zepdrix):

Since we're given an interval, we have to also check the `end points`. So if you've found the critical points correctly. Plug them into the function separately, then also plug in the end points (0 and 2 for part a). Then simply compare all of your values that they produce. The largest will be your max, smallest your min.

zepdrix (zepdrix):

f(0) = ? f(-1) = ? f(1) = ? f(2) = ?

OpenStudy (anonymous):

okay so the critical points are +/- 1

zepdrix (zepdrix):

ok sounds good :)

OpenStudy (anonymous):

so now i plug them into the problem right?

zepdrix (zepdrix):

Yes, into the original function (what it looked like before you took a derivative).

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

(1)^3 - 3(1) + 1 = -1

OpenStudy (anonymous):

so -1 and -3

zepdrix (zepdrix):

f(1) = -1 f(-1) = -3 Ok good. We also need to check the end points now. The ends of our interval. Plug 0 and 2 into the function. f(0) = ? f(2) = ?

zepdrix (zepdrix):

Woops I think f(-1) should be giving us +3, not -3.

OpenStudy (anonymous):

yes you're right its positive 3, so for f(0) = -2 and f(2) its 3

zepdrix (zepdrix):

f(0) = 0^3 - 3(0) + 1 f(0) = 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so -1, 3, 1 and 3

zepdrix (zepdrix):

Ok so we now have the function value at our critical points and at the ends of our interval. Our interval was from 0 to 2. Hmmm -1 is NOT in that interval is it?

OpenStudy (anonymous):

no

zepdrix (zepdrix):

So we'll throw that point out. These are the points we care about. f(0) = 1 f(1) = -1 f(2) = 3 Which one is the largest? That will be our maximum.

OpenStudy (anonymous):

ok 3 is the largest

zepdrix (zepdrix):

Ok good! We could write that like this perhaps, as a coordinate pair. Max: (2,3) How about your minimum? :)

OpenStudy (anonymous):

the minimum is -1

OpenStudy (anonymous):

Min : (1,-1)

zepdrix (zepdrix):

Yes very good. When they ask for `critical points` they simply want the X value. When they ask for `max` or `min points` they want it as an ordered pair usually.

OpenStudy (anonymous):

ok thanks